# Current address: ]College of William and Mary, Williamsburg, Virginia 23187, USA Current address: ]Thomas Jefferson National Accelerator Facility, Newport News, Virginia 23606, USA

### v3

@@ VI Results on the Structure Functions
We present in this section results on the proton and the deuteron structure functions and , which were reported previously in Refs. [21, 22]. We also present new results on the neutron as extracted from the proton and deuteron data, as well as their isovector combination .
@@ VI.1 Results on and for the Proton and Deuteron
Our final results for and of the proton and the deuteron are shown in Figs. 10 through 13. Our results are consistent with the previous data [103, 102] where their coverage overlap, but extend the measured range to three times smaller values, below the pion mass squared (). This makes it possible to rigorously test EFT calculations for the nucleon spin structure functions. Comparing to our existing model, which was based on existing data prior to EG4, one can see that it tends to be more positive than the data points, especially at lower . Our new data can help improve this model.
@@ VI.2 Results on neutron
From our results on of the proton and the deuteron, we further extracted values for the neutron spin structure function, , which were not available previously. This extraction is based on the simplified model of the deuteron structure functions described in Section II.5, where the structure functions of the deuteron are expressed as sum of the two nucleon structure functions, convoluted with their spin and momentum distributions in the deuteron [115]. To extract the deconvoluted structure functions for a neutron at rest, we followed a similar procedure as for the extraction of from the data on the proton and deuteron, described in detail below.
As a first step, we took our results and converted them to the contribution from the bound proton to the deuteron spin structure function using the folding technique of Ref. [115]. The bound were then subtracted from our results to obtain data for the bound neutron . The data were compared with expected values for the bound neutron, , produced using our model of the free neutron processed by the same folding technique, and any observed difference is attributed to how the input model of the free neutron differs from reality. We then produced a modification of the base model where was changed by () in the input, and the difference between the modified and the original model provides information on how much the model of the free neutron should be adjusted to match data. More specifically, for the free neutron is extracted as:
@@
Finally, we show the isovector combination in Fig. 15. Comparing with Figs.10, 11, and 14, one clearly observes the suppression in the isovector component of the . This has important consequences for the Bjorken integral (i.e., the first moment of ), see Section VII.6, and its comparison to EFT [151]. Like the other EG4 results on and , Fig. 15 also makes clear that in the domain covered by EG4, the parameterization of the world data (continuous blue line) needs to be revised.
@@ VII Results on Moments of Structure Functions
Similar to the previous section(s), we summarize the procedure to form the moments of the proton and the deuteron structure functions, and present the results that were reported previously in Refs. [21, 22] but with updated PT calculations. We also present new results on the moments of the neutron extracted directly from the moments of the proton and the deuteron, as well as their isospin combinations and .
@@ VII.2 Latest EFT Moment Predictions
In the following sections, we will compare the EG4 result on the moments to theoretical predictions, in particular from EFT. We will display only two recent predictions, from Ref. [59] and Ref. [60], as they supersede earlier EFT calculations [152, 79, 81, 83]. The crucial advance of the latest calculations is that they account rigorously for the excitation. At the same time, the approaches by Refs. [59] and [60] have significant differences. The chief reason is the effective accounting of higher-order contributions of the EFT series by Ref. [60], which uses a phenomenological form factor. It softens the steep -dependence otherwise seen in Ref. [59]. The softening form factor of Ref. [60] accounts for their typically better description of the -dependence of moments. The methodological differences between the two calculations are detailed in Section VII.7.
@@ VII.3 Moments of the Proton Spin Structure functions
They agree with the most recent EFT predictions [59, 60] up to GeV2, above which only that of Ref. [60] agrees with the data. The new results generally agree with a previous experiment [102] in the overlapping region within their uncertainties. However, it is visible that the parameterization based on Ref. [102] (purple solid curve) can be improved by including the new results. The phenomenological models Burkert and Ioffe [53, 54] and Pasechnik et al. [56] agree well with the new results for all values. They predict the behavior of in the non-perturbative domain of QCD and beyond: The Burkert and Ioffe model [53, 54] extrapolates DIS data using vector meson dominance and a parameterization of resonance contributions; while the Pasechnik et al. model [56], improved upon the earlier Soffer-Teryaev model [55], employs the analytical perturbation theory (APT) approach that extrapolates DIS data to low , accounting for the mild -dependence of . Lastly, the MAID unitary isobar phenomenological model [113] is a parametrization of experimental pion production data based on a partial wave analysis in the resonance region, and includes nonresonant background. MAID does not account for multi-particle final states and therefore provides a partial in the resonance region, without the low- (high-) contribution. Consequently, it can be roughly compared with the experimentally measured (open blue circles), although it does not necessarily have the same integration limits. With these caveats in mind, MAID and the EG4 measured sum agree well up to GeV2, after which the model is below the data.
@@
Our results on (Eq. (30)) are shown in Fig. 17. Compared with the EFT predictions, similar conclusions as for are reached. The MAID model still agrees with the experimentally measured integral within uncertainties. Finally, the difference between the new results and the parameterization of previous data [102] is more visible and the parameterization can clearly be improved.
Extrapolating our results to yields [22]
@@
This agrees well with the GDH sum rule prediction and the experimental photoproduction result (stat) (syst) [65, 67, 68, 66]. We note that the uncertainties from Refs. [65, 67, 68, 66] and Eq. (57) are comparable. For EG4, the uncertainty due to the extrapolation is compensated by the fact that inclusive electroproduction automatically sums over all reaction channels, removing uncertainties due to the detection of all final states needed in photoproduction. The EG4 data thus provide the first test of the GDH sum rule with a different technique than photoproduction.
@@
One more moment that can be formed from the EG4 data is the generalized longitudinal spin polarizability (Eq. (35)). Compared to (Eq. (30)), the -weighting in Eq. (35) indicates that and have different contributions from both systematic and model uncertainties. Results for are shown in Fig. 18. They agree with EFT calculation from Bernard et al. [59] only at very low (below GeV2) and agree marginally with Alarcon et al. [60], though both calculations have large uncertainties. On the other hand, when combined with the photoproduction data point [153], the EG4 results support the larger slope at very low predicted by Bernard et al. [59]. The agreement between the measured integrals and the MAID model is similar to both and . The recent calculation by Bigazzi et al. [154] (not shown in the figure) based on the anti-de Sitter/conformal field theory (AdS/CFT) correspondence [155] applied to QCD [156] recovers the sign and qualitative behavior of , albeit up to a normalization factor.
@@ VII.4 Moments of the Deuteron Spin Structure Functions
Similar observations as for the proton measurements can be drawn: the EG4 data agree with the previous data from CLAS [102] (and from SLAC E143 [107], not shown in the figure). The measured and agree with the EFT results for the sum of proton and neutron (multiplied with the D-state depolarization, see Section II.5) of Alarcón et al. [60], and of Bernard et al. [59] for the lowest points. The models of Pasechnik et al. [56] and Burkert-Ioffe [53, 54] agree well with the data, while the parameterization [102] is systematically slightly larger. The comparison between the MAID model and the data is similar to that of the proton for all three integrals.
Extrapolating the data to yields [21]
which agrees at the level with the GDH sum rule prediction , obtained using Eq. (37), where the uncertainty is from the deuteron D-state correction. This can be compared with the data using real photons of energy between GeV, (stat) (syst) [65]. Note that the systematic uncertainty of [65] does not include that from the unmeasured low and large contributions.
@@ VII.5 Moments of the Neutron Spin Structure Functions
The resulting and are shown in Figs. 22 and 23, respectively. They agree with previous data from E143 [107], EG1a [94, 95] and EG1b [102, 103] where the neutron information was extracted from the combined deuteron and proton data, as well as those from E94-010 [93] and E97-110 [71] for which the neutron information was extracted from 3He data. There is also agreement with the phenomenological models of Burkert and Ioffe [53, 54] and Pasechnik et al. [56] (when available) and with the EFT predictions from Bernard et al. [59] (up to GeV2) and, to a lesser extent, with Alarcón et al. [60]. The MAID model prediction is systematically more positive than the data. Like in the case of , the calculation of [154] (not shown in the figure) agrees in sign and behavior with the , up to a similar normalization factor as for .
@@
Results on the generalized longitudinal spin polarizability on the neutron are shown in Fig. 24, along with earlier results from JLab EG1b [102, 103]), E97-110 [71] and E94-010 [93]). The EG4 neutron results generally agree with the earlier data, though there is a tension between EG4 and E97-110 for GeV2. Adding the EG4 deuteron and proton systematic uncertainties linearly, rather than in quadrature, does not remove the tension. It contrasts with the agreement between EG4 and E97-110 for the neutron data on or , suggesting the discrepancy arises from the high- (near pion threshold) contribution. The EG4 results agree with EFT calculation from Ref. [59] up to GeV2 but not with that of Alarcón et al. [60].
@@ VII.6 Isospin Analysis on the Moments of Spin Structure Functions
where is given under Eq. (38) and with the deuteron moments understood as “per nucleus”. Alternatively to using the EG4 proton and deuteron data, one can use the proton results from EG4 and the neutron results obtained from 3He data such as those from JLab E97-110 [71] to perform the isospin separation.
The isovector moments are expected to be simpler to compute than the isoscalar or nucleon moments. For lattice QCD simulations, this is because the disconnected diagrams are suppressed [17]. For EFT calculations, e.g., Refs. [151, 60], the inclusion of the excitation is often difficult, which would cancel in the isovector combination. However, the consequence of the expectation that the EFT predictions are more robust has been upheld only with data on the Bjorken sum [42, 43, 44], but not with [43], nor with the longitudinal-transverse interference polarizability [19, 104] for which the contribution is also expected to cancel. Our results provide improved tests of this expectation thanks to the lower coverage and higher precision of EG4 data.
@@ VII.6.1 Results on the Bjorken Sum
The isovector combination is the Bjorken sum [27, 28], see Eqs. (25)-(32). Our results on are shown in Fig. 25, compared with data from earlier experiments [106, 42, 43, 44]. Besides the models already discussed in connection with Figs. 16-24, we also show predictions from holographic light-front QCD (HLFQCD) [156]. In HLFQCD, the QCD coupling is derived in the scheme following the effective charge prescription [157, 7, 8], from which the Bjorken sum can be obtained [158, 159, 160].
@@
Note the comparison between the results obtained from EG4’s proton and deuteron data and those extracted from the EG4 proton and E97-110 neutron (3He) data. An agreement between the two different approaches was not assured since the nuclear structures of the deuteron and 3He are quite different: 3He is strongly bound compared to the deuteron, therefore more sensitive to potential nuclear modification of the proton. Our treatment of the deuteron is minimal and consists of accounting for its D-state [116, 117, 118, 119, 120], while treatment of the 3He data accounted for the nucleon effective polarizations [162]. As shown in Fig. 25, the two approaches agree well with each other. This suggests that, at least for , the nuclear corrections applied to the deuteron and 3He data to obtain the neutron information appear sufficient at these low values.
@@
Both the EG4 and EG4/E97-110 results on are systematically above EFT and model predictions, except that of HLFQCD [58] and [161], which agree with the data. To allow a quantitative comparison between the data and theory, we employed a fit of the form to the entire world data set, based on Eq. (32) [52], where is expected to be the GDH slope. Our results are shown in Table 4. The “” uncertainty designates the point-to-point uncorrelated uncertainty. It is the quadratic sum of the statistical uncertainty and a fraction of the systematic uncertainty determined so that /n.d.f = 1 for the best fit. The “” uncertainty is the correlated uncertainty estimated from the remaining fraction of the systematic uncertainty.
@@
We note that our value for deviates significantly from that expected from the GDH sum rule. This is partially due to the large cancellation for the GDH sum rule in the combination, which is only 1/4 of the value for the individual nucleons. Experimentally, because of the nearly perfect cancellation of the large contribution from the resonance in the difference, the isovector integral is much more strongly affected by small- contributions than the other moments, with relatively large uncertainties for . Of course, our results could also indicate a much stronger –dependence near the photon point than can be captured by a 2-parameter fit. In any case, they do not indicate a violation of the GDH sum rule since, as shown in the previous section, our results for the proton and neutron are individually consistent with the sum rule within (1 - 1.3). Further discussions on the possible discrepancy between the parameter and the GDH slope can be found in Ref. [52].
@@
| GDH sum rule [61] | 0.0618 | - | ||||
| EFT Bernard et al. [59] | 0.07 | 0.3 | ||||
| EFT Alarcón et al. [60] | 0.066(4) | 0.25(12) | ||||
| Burkert-Ioffe [53, 54] | 0.09 | 0.3 | ||||
| Pasechnik et al. [56] | 0.09 | 0.4 | ||||
| HLFQCD [58] | 0.177 | -0.067 |

### v4

@@ VI Results on the Structure Functions
We present in this section results on the proton and the deuteron structure functions and , which were reported previously in Refs. [CLAS:2017ozc, CLAS:2021apd]. We also present new results on the neutron as extracted from the proton and deuteron data, as well as their isovector combination .
@@ VI.1 Results on and for the Proton and Deuteron
Our final results for and of the proton and the deuteron are shown in Figs. 10 through 13. Our results are consistent with the previous data [CLAS:2015otq, CLAS:2017qga] where their coverage overlap, but extend the measured range to three times smaller values, below the pion mass squared (). This makes it possible to rigorously test EFT calculations for the nucleon spin structure functions. Comparing to our existing model, which was based on existing data prior to EG4, one can see that it tends to be more positive than the data points, especially at lower . Our new data can help improve this model.
@@ VI.2 Results on neutron
From our results on of the proton and the deuteron, we further extracted values for the neutron spin structure function, , which were not available previously. This extraction is based on the simplified model of the deuteron structure functions described in Section II.5, where the structure functions of the deuteron are expressed as sum of the two nucleon structure functions, convoluted with their spin and momentum distributions in the deuteron [Kahn:2008nq]. To extract the deconvoluted structure functions for a neutron at rest, we followed a similar procedure as for the extraction of from the data on the proton and deuteron, described in detail below.
As a first step, we took our results and converted them to the contribution from the bound proton to the deuteron spin structure function using the folding technique of Ref. [Kahn:2008nq]. The bound were then subtracted from our results to obtain data for the bound neutron . The data were compared with expected values for the bound neutron, , produced using our model of the free neutron processed by the same folding technique, and any observed difference is attributed to how the input model of the free neutron differs from reality. We then produced a modification of the base model where was changed by () in the input, and the difference between the modified and the original model provides information on how much the model of the free neutron should be adjusted to match data. More specifically, for the free neutron is extracted as:
@@
Finally, we show the isovector combination in Fig. 15. Comparing with Figs.10, 11, and 14, one clearly observes the suppression in the isovector component of the . This has important consequences for the Bjorken integral (i.e., the first moment of ), see Section VII.6, and its comparison to EFT [Burkert:2000qm]. Like the other EG4 results on and , Fig. 15 also makes clear that in the domain covered by EG4, the parameterization of the world data (continuous blue line) needs to be revised.
@@ VII Results on Moments of Structure Functions
Similar to the previous section(s), we summarize the procedure to form the moments of the proton and the deuteron structure functions, and present the results that were reported previously in Refs. [CLAS:2017ozc, CLAS:2021apd] but with updated PT calculations. We also present new results on the moments of the neutron extracted directly from the moments of the proton and the deuteron, as well as their isospin combinations and .
@@ VII.2 Latest EFT Moment Predictions
In the following sections, we will compare the EG4 result on the moments to theoretical predictions, in particular from EFT. We will display only two recent predictions, from Ref. [Bernard:2012hb] and Ref. [Alarcon:2020icz], as they supersede earlier EFT calculations [Bernard:1992nz, Ji:1999pd, Bernard:2002bs, Kao:2002cp]. The crucial advance of the latest calculations is that they account rigorously for the excitation. At the same time, the approaches by Refs. [Bernard:2012hb] and [Alarcon:2020icz] have significant differences. The chief reason is the effective accounting of higher-order contributions of the EFT series by Ref. [Alarcon:2020icz], which uses a phenomenological form factor. It softens the steep -dependence otherwise seen in Ref. [Bernard:2012hb]. The softening form factor of Ref. [Alarcon:2020icz] accounts for their typically better description of the -dependence of moments. The methodological differences between the two calculations are detailed in Section VII.7.
@@ VII.3 Moments of the Proton Spin Structure functions
They agree with the most recent EFT predictions [Bernard:2012hb, Alarcon:2020icz] up to GeV2, above which only that of Ref. [Alarcon:2020icz] agrees with the data. The new results generally agree with a previous experiment [CLAS:2017qga] in the overlapping region within their uncertainties. However, it is visible that the parameterization based on Ref. [CLAS:2017qga] (purple solid curve) can be improved by including the new results. The phenomenological models Burkert and Ioffe [Burkert:1992tg, Burkert:1993ya] and Pasechnik et al. [Pasechnik:2010fg] agree well with the new results for all values. They predict the behavior of in the non-perturbative domain of QCD and beyond: The Burkert and Ioffe model [Burkert:1992tg, Burkert:1993ya] extrapolates DIS data using vector meson dominance and a parameterization of resonance contributions; while the Pasechnik et al. model [Pasechnik:2010fg], improved upon the earlier Soffer-Teryaev model [Soffer:2004ip], employs the analytical perturbation theory (APT) approach that extrapolates DIS data to low , accounting for the mild -dependence of . Lastly, the MAID unitary isobar phenomenological model [Drechsel:1998hk] is a parametrization of experimental pion production data based on a partial wave analysis in the resonance region, and includes nonresonant background. MAID does not account for multi-particle final states and therefore provides a partial in the resonance region, without the low- (high-) contribution. Consequently, it can be roughly compared with the experimentally measured (open blue circles), although it does not necessarily have the same integration limits. With these caveats in mind, MAID and the EG4 measured sum agree well up to GeV2, after which the model is below the data.
@@
Our results on (Eq. (30)) are shown in Fig. 17. Compared with the EFT predictions, similar conclusions as for are reached. The MAID model still agrees with the experimentally measured integral within uncertainties. Finally, the difference between the new results and the parameterization of previous data [CLAS:2017qga] is more visible and the parameterization can clearly be improved.
Extrapolating our results to yields [CLAS:2021apd]
@@
This agrees well with the GDH sum rule prediction and the experimental photoproduction result (stat) (syst) [GDH:2001zzk, GDH:2003xhc, GDH:2005noz, Ahrens:2006yx]. We note that the uncertainties from Refs. [GDH:2001zzk, GDH:2003xhc, GDH:2005noz, Ahrens:2006yx] and Eq. (57) are comparable. For EG4, the uncertainty due to the extrapolation is compensated by the fact that inclusive electroproduction automatically sums over all reaction channels, removing uncertainties due to the detection of all final states needed in photoproduction. The EG4 data thus provide the first test of the GDH sum rule with a different technique than photoproduction.
@@
One more moment that can be formed from the EG4 data is the generalized longitudinal spin polarizability (Eq. (35)). Compared to (Eq. (30)), the -weighting in Eq. (35) indicates that and have different contributions from both systematic and model uncertainties. Results for are shown in Fig. 18. They agree with EFT calculation from Bernard et al. [Bernard:2012hb] only at very low (below GeV2) and agree marginally with Alarcon et al. [Alarcon:2020icz], though both calculations have large uncertainties. On the other hand, when combined with the photoproduction data point [Gurevich:2017cpt], the EG4 results support the larger slope at very low predicted by Bernard et al. [Bernard:2012hb]. The agreement between the measured integrals and the MAID model is similar to both and . The recent calculation by Bigazzi et al. [Bigazzi:2023odl] (not shown in the figure) based on the anti-de Sitter/conformal field theory (AdS/CFT) correspondence [Maldacena:1997re] applied to QCD [Brodsky:2014yha] recovers the sign and qualitative behavior of , albeit up to a normalization factor.
@@ VII.4 Moments of the Deuteron Spin Structure Functions
Similar observations as for the proton measurements can be drawn: the EG4 data agree with the previous data from CLAS [CLAS:2017qga] (and from SLAC E143 [E143:1996vck], not shown in the figure). The measured and agree with the EFT results for the sum of proton and neutron (multiplied with the D-state depolarization, see Section II.5) of Alarcón et al. [Alarcon:2020icz], and of Bernard et al. [Bernard:2012hb] for the lowest points. The models of Pasechnik et al. [Pasechnik:2010fg] and Burkert-Ioffe [Burkert:1992tg, Burkert:1993ya] agree well with the data, while the parameterization [CLAS:2017qga] is systematically slightly larger. The comparison between the MAID model and the data is similar to that of the proton for all three integrals.
Extrapolating the data to yields [CLAS:2017ozc]
which agrees at the level with the GDH sum rule prediction , obtained using Eq. (37), where the uncertainty is from the deuteron D-state correction. This can be compared with the data using real photons of energy between GeV, (stat) (syst) [GDH:2001zzk]. Note that the systematic uncertainty of [GDH:2001zzk] does not include that from the unmeasured low and large contributions.
Regarding the generalized spin polarizability , the EG4 results are clearly outside the range predicted by Alarcón et al. [Alarcon:2020icz], but agree with the Bernard et al. calculation for the lowest points. The parameterization [CLAS:2017qga] does not describe the data well below GeV2, similarly to the case.
@@ VII.5 Moments of the Neutron Spin Structure Functions
The resulting and are shown in Figs. 22 and 23, respectively. They agree with previous data from E143 [E143:1996vck], EG1a [CLAS:2002knl, CLAS:2003rjt] and EG1b [CLAS:2017qga, CLAS:2015otq] where the neutron information was extracted from the combined deuteron and proton data, as well as those from E94-010 [JeffersonLabE94-010:2003dvv] and E97-110 [JeffersonLabE97-110:2019fsc] for which the neutron information was extracted from 3He data. There is also agreement with the phenomenological models of Burkert and Ioffe [Burkert:1992tg, Burkert:1993ya] and Pasechnik et al. [Pasechnik:2010fg] (when available) and with the EFT predictions from Bernard et al. [Bernard:2012hb] (up to GeV2) and, to a lesser extent, with Alarcón et al. [Alarcon:2020icz]. The MAID model prediction is systematically more positive than the data. Like in the case of , the calculation of [Bigazzi:2023odl] (not shown in the figure) agrees in sign and behavior with the , up to a similar normalization factor as for .
@@
Results on the generalized longitudinal spin polarizability on the neutron are shown in Fig. 24, along with earlier results from JLab EG1b [CLAS:2017qga, CLAS:2015otq]), E97-110 [JeffersonLabE97-110:2019fsc] and E94-010 [JeffersonLabE94-010:2003dvv]). The EG4 neutron results generally agree with the earlier data, though there is a tension between EG4 and E97-110 for GeV2. Adding the EG4 deuteron and proton systematic uncertainties linearly, rather than in quadrature, does not remove the tension. It contrasts with the agreement between EG4 and E97-110 for the neutron data on or , suggesting the discrepancy arises from the high- (near pion threshold) contribution. The EG4 results agree with EFT calculation from Ref. [Bernard:2012hb] up to GeV2 but not with that of Alarcón et al. [Alarcon:2020icz].
@@ VII.6 Isospin Analysis on the Moments of Spin Structure Functions
where is given under Eq. (38) and with the deuteron moments understood as “per nucleus”. Alternatively to using the EG4 proton and deuteron data, one can use the proton results from EG4 and the neutron results obtained from 3He data such as those from JLab E97-110 [JeffersonLabE97-110:2019fsc] to perform the isospin separation.
The isovector moments are expected to be simpler to compute than the isoscalar or nucleon moments. For lattice QCD simulations, this is because the disconnected diagrams are suppressed [Deur:2018roz]. For EFT calculations, e.g., Refs. [Burkert:2000qm, Alarcon:2020icz], the inclusion of the excitation is often difficult, which would cancel in the isovector combination. However, the consequence of the expectation that the EFT predictions are more robust has been upheld only with data on the Bjorken sum [Deur:2004ti, Deur:2008ej, Deur:2014vea], but not with [Deur:2008ej], nor with the longitudinal-transverse interference polarizability [JeffersonLabE94010:2004ekh, E97-110:2021mxm] for which the contribution is also expected to cancel. Our results provide improved tests of this expectation thanks to the lower coverage and higher precision of EG4 data.
@@ VII.6.1 Results on the Bjorken Sum
The isovector combination is the Bjorken sum [Bjorken:1966jh, Bjorken:1969mm], see Eqs. (25)-(32). Our results on are shown in Fig. 25, compared with data from earlier experiments [E143:1994vcg, Deur:2004ti, Deur:2008ej, Deur:2014vea]. Besides the models already discussed in connection with Figs. 16-24, we also show predictions from holographic light-front QCD (HLFQCD) [Brodsky:2014yha]. In HLFQCD, the QCD coupling is derived in the scheme following the effective charge prescription [Grunberg:1980ja, Deur:2016tte, Deur:2023dzc], from which the Bjorken sum can be obtained [Deur:2005cf, Deur:2008rf, Deur:2022msf].
@@
Note the comparison between the results obtained from EG4’s proton and deuteron data and those extracted from the EG4 proton and E97-110 neutron (3He) data. An agreement between the two different approaches was not assured since the nuclear structures of the deuteron and 3He are quite different: 3He is strongly bound compared to the deuteron, therefore more sensitive to potential nuclear modification of the proton. Our treatment of the deuteron is minimal and consists of accounting for its D-state [Lacombe:1980dr, Machleidt:1987hj, Zuilhof:1980ae, Kotthoff:1976fg, Desplanques:1988mp], while treatment of the 3He data accounted for the nucleon effective polarizations [CiofidegliAtti:1996cg]. As shown in Fig. 25, the two approaches agree well with each other. This suggests that, at least for , the nuclear corrections applied to the deuteron and 3He data to obtain the neutron information appear sufficient at these low values.
@@
Both the EG4 and EG4/E97-110 results on are systematically above EFT and model predictions, except that of HLFQCD [Brodsky:2010ur] and [Gabdrakhmanov:2023rjt], which agree with the data. To allow a quantitative comparison between the data and theory, we employed a fit of the form to the entire world data set, based on Eq. (32) [Deur:2021klh], where is expected to be the GDH slope. Our results are shown in Table 4. The “” uncertainty designates the point-to-point uncorrelated uncertainty. It is the quadratic sum of the statistical uncertainty and a fraction of the systematic uncertainty determined so that /n.d.f = 1 for the best fit. The “” uncertainty is the correlated uncertainty estimated from the remaining fraction of the systematic uncertainty.
@@
We note that our value for deviates significantly from that expected from the GDH sum rule. This is partially due to the large cancellation for the GDH sum rule in the combination, which is only 1/4 of the value for the individual nucleons. Experimentally, because of the nearly perfect cancellation of the large contribution from the resonance in the difference, the isovector integral is much more strongly affected by small- contributions than the other moments, with relatively large uncertainties for . Of course, our results could also indicate a much stronger –dependence near the photon point than can be captured by a 2-parameter fit. In any case, they do not indicate a violation of the GDH sum rule since, as shown in the previous section, our results for the proton and neutron are individually consistent with the sum rule within (1 - 1.3). Further discussions on the possible discrepancy between the parameter and the GDH slope can be found in Ref. [Deur:2021klh].
@@
| GDH sum rule [Gerasimov:1965et] | 0.0618 | - | ||||
| EFT Bernard et al. [Bernard:2012hb] | 0.07 | 0.3 | ||||
| EFT Alarcón et al. [Alarcon:2020icz] | 0.066(4) | 0.25(12) | ||||
| Burkert-Ioffe [Burkert:1992tg, Burkert:1993ya] | 0.09 | 0.3 | ||||
| Pasechnik et al. [Pasechnik:2010fg] | 0.09 | 0.4 | ||||
| HLFQCD [Brodsky:2010ur] | 0.177 | -0.067 |